MEDICATION CALCULATION #70

Order: Leukeran 4.5 mg/m(squared)/single dose. Child's weight: 92 pounds; height is 58 inches. Use the nomogram below to determine body surface area (m(squared)). Child's drug dosage: mg/kg/day in divided doses. Available: Leukeran (chlorambucil) 2 mg tablet

a. What is the child' body surface area?

b. How many tablet(s) should the child receive?

Please show your step by step calculations.
Answer #1

A, Body surface area= ( height (cm)X wt (kg)/3600)

92pounds = 41.7(42kg)

58inches = 147cm

147 X 42/3600= 1.7

BSA = 1.7

B, Available dose = 2mg

Ordered dose= 4.5mg

= Ordered dose/ available dose

4.5/2 = 2.25tablets the child should receive..